# IA2: Alignment with ICL Activations Improves Supervised Fine-Tuning

### v1

@@ Abstract
Supervised Fine-Tuning (SFT) is used to specialize model behavior by training weights to produce intended target responses for queries. In contrast, In-Context Learning (ICL) adapts models during inference with instructions or demonstrations in the prompt. ICL can offer better generalizability and more calibrated responses compared to SFT in data scarce settings, at the cost of more inference compute. In this work, we ask the question: Can ICL’s internal computations be used to improve the qualities of SFT? We first show that ICL and SFT produce distinct activation patterns, indicating that the two methods achieve adaptation through different functional mechanisms. Motivated by this observation and to use ICL’s rich functionality, we introduce ICL Activation Alignment (IA2), a self-distillation technique which aims to replicate ICL’s activation patterns in SFT models and incentivizes ICL-like internal reasoning. Performing IA2 as a priming step before SFT significantly improves the accuracy and calibration of model outputs, as shown by our extensive empirical results on 12 popular benchmarks and 2 model families. This finding is not only practically useful, but also offers a conceptual window into the inner mechanics of model adaptation.
@@ 5 Experimental Results
Adaptation Tasks: In the single-token setting, the output $Y_{i}$ for each sample is a single token, typically a category/class label, True/False, or MCQ choice. In the multi-token setting, the output $Y_{i}$ can be multiple tokens long and variable in length. This is the general case when models are used for open-ended generation. For both settings, we test and compare our proposed SFT pipeline on various tasks using multiple models. In many cases, we also test the adaptation methods on OOD datasets to study their behavior under distribution shift. We enlist the tasks under each category below:
@@
In the multi-token setting, the model outputs a sequence of reasoning tokens (not inside the <<think>> tags, just standard text) before the answer in a specific format illustrated in the demonstrations. The actual answer needs to be extracted from the generated text to evaluate the model performance.
@@
Data Setup: As our focus is on the data-scarce few-shot setting where ICL is typically used, we create multiple training datasets $\mathcal{D}_{\mathcal{T}}=\{(X_{i},Y_{i})\}_{i=1}^{N}$ for each task by varying $N$ in orders of 2:[2,4,8,16,]2:[2,4,8,16,...]. We create 5 different sets at each $N$ value to average out outlier effects on our final quantitative analysis. Importantly, we collect ICL activation tensors $A^{i}_{\text{ICL}}\hskip 2.0pt\forall X_{i}$ for IA2 training by using the remaining N1N-1 samples as ICL demonstrations (in random order), effectively reusing the same samples each time. This makes sure that we use the same data in all adaptation methods. For evaluation, we sampled a different set of 500 samples for each task. Additional details about the datasets and how they are processed can be found in §A.
@@
Training Setup: We use two model families to test each training method across all tasks. We use the Qwen3-4B-Base model on every task (single/multi-token). In addition, we also use the Llama-3.2-1B model on every single-token task except SciQr for which we use Llama-3.2-3B. This is because the 1B Llama model was unable to improve above random baseline for any $N$ upto 128. For the multi-token tasks, we use the Llama-3.2-1B-Instruct model as the secondary model. For training models, we use LoRAs with rank 8 to modify the $W_{Q},W_{K}$ and $W_{O}$ matrices of self-attention layers. We chose these models and weight combinations reasonably on the basis of available resources, as we trained more than 13,000 models in total for our experiments. Training in multi-token setting is much more resource extensive. As each sample has a long answer, we decided to only use $N$ upto 16 for Llama and $N$ upto 8 for Qwen models in multi-token setting. We also fix the maximum generated tokens ($G=200$) during training for efficiency. This is used in IA2 for collecting the activation tensor $\hat{A}_{i}\in\mathbb{R}^{L\times 200\times d}\hskip 2.0pt\forall x_{i}$. For SFT, $G$ is equal to the number of tokens in the ground truth answer. Each (method, dataset) combination is trained with 3 learning rates (slow:1e-4, medium:3e-4, fast:1e-3) and the best performing one is chosen to neutralize the effect of hyperparameter selection on method performance. Additional details in §B.
@@ 6 Discussion
IA2 + SFT: One natural alternative to IA2 → SFT training is IA2 + SFT, i.e., using both $\mathcal{L}_{\text{{{IA2}}}}$ and $\mathcal{L}_{\text{SFT}}$ at the same time rather than sequentially. This approach poses a practical challenge. $\hat{A}$ collected as the target activation tensor for IA2 uses the model’s own generated response tokens for any $G>1$. We call it the ICL response, which could be very different (even in length) from the ground truth response in the dataset. This makes the two objectives incompatible to train together. However, we could perform self-distillation using ICL responses, similar to Snell et al. (2022) but including an additional IA2 signal. We performed this training with a unified loss $\mathcal{L}_{\text{{{IA2}}}}+\beta\cdot\mathcal{L}_{\text{SFT}}$ and varied $\beta$ with 4 values spanning the spectrum between IA2 only and SFT only training smoothly (details in §B). After choosing the best performing (learning rate, $\beta$) combination across 5 random seeds, we report the performance for multi-token experiments in Table 3. Note that the performance of SFT only and IA2 → SFT training drops when compared to training on ground truth responses (from Table 2), but IA2 + SFT extracts significantly more performance out of ICL responses when compared to SFT only training. This highlights the richness of the IA2 training signal. Additional results of IA2 + SFT training can be found in §C.
@@ 7 Related Work
Context Distillation: A large body of work targets to improve LLM adaptation using some form of compression of explicit context signals like us. These range from using intermediate contemplation tokens for compression (Cheng & Van Durme, 2024), internalizing context through generating adapters on the fly (Chen et al., 2025; Charakorn et al., 2025), using related text to enhance distillation (Zhu et al., 2025; Choi et al., 2025), or compressing context directly into weights (Shen et al., 2025; Deng et al., 2024; Yu et al., 2024; Snell et al., 2022; Chen et al., 2024b; Shin et al., 2024). These works aim to distill the effect of given contexts on model outputs using token response signals. Our work is complementary to these works and improves the distillation effects through the model’s own ICL processing signals, an idea with supporting prior work (Aguilar et al., 2020; Jin et al., 2024; Yang et al., 2024).
@@ 8 Conclusion and Future Work
In this work, we used model activations to probe the functional behavior of LLMs and found a key distinction: ICL and SFT achieve adaptation through different mechanisms, and ICL often encodes richer, more generalizable patterns. Motivated by this, we introduced IA2, an SFT priming technique that enables ICL-like internal behavior in the model. This simple step shifts models into a more adaptable weight subspace, which is inaccessible to SFT alone. Extensive experiments show that IA2 consistently improves the accuracy and calibration of SFT. In the future, we aim to test other effects like impact on diversity and catastrophic forgetting of IA2 based SFT to measure how IA2 can help in post-training LLMs.
@@ Appendix A Additional Dataset Details
Parsing multi-token answers: Different datasets had different method for parsing the answer from response tokens.
  • For GSM8K and GSM8Ks, the ground truth answers had the following pattern: “<<reasoning steps>> #### <<numerical answer>>”. We do not evaluate the correctness or conciseness of the reasoning steps and only parse the numerical answer out from the response and match it exactly with the ground truth answer for calculating accuracy. We also use other common answer parsing techniques like matching with“The answer is <<value>>”, to allow for slight variations in output format. Our exact parsing function can be seen in our code.
@@
We did not use publically avaliable evaluation frameworks like lm-evaluation-harness on known benchmarks like GSM8K because we wanted more fine-grained control over which samples were used in the evaluation process. As our parsing functions remain consistent across all training/adaptation methods, the relative performance between methods still gives a good estimate of how they would perform on standard evaluations.
@@ Appendix B Additional Experimental Details
LoRA details: We used LoRA rank =8=8, $\alpha=8$ and target modules are attention weights $W_{Q},W_{K}$ and $W_{O}$ for all our experiments. We settled on these parameters on the basis of generally used values for small targeted adaptations, as additional explorations require substantial compute resources for the scale of our experiments. The target modules capture the cross-attention parts ($Q,K$) and final projection $O$ to allow for better adaptability. However, these selections are subject to improvement with more exploration.
@@
Subspace Overlap: The subspace overlap plot (Figure 4) uses models from all single and multi token experiments. The high spread of subspace overlap between IA2 only and IA2 → SFT models is because in many cases, the IA2 model itself was already output aligned, so the secondary SFT training did not change the model much before convergence.
[…]

### v2

@@ Abstract
Supervised Fine-Tuning (SFT) is used to specialize model behavior by training weights to produce intended target responses for queries. In contrast, In-Context Learning (ICL) adapts models during inference with instructions or demonstrations in the prompt. ICL can offer better generalizability and more calibrated responses compared to SFT in data scarce settings, at the cost of more inference compute. In this work, we ask the question: Can ICL’s internal computations be used to improve the qualities of SFT? We first show that ICL and SFT produce distinct activation patterns, indicating that the two methods achieve adaptation through different functional mechanisms. Motivated by this observation and to use ICL’s rich functionality, we introduce ICL Activation Alignment (IA2), a self-distillation technique which aims to replicate ICL’s activation patterns in SFT models and incentivizes ICL-like internal reasoning. Performing IA2 as a priming step before SFT significantly improves the accuracy and calibration of model outputs, as shown by our extensive empirical results on 12 popular benchmarks and two model families. This finding is not only practically useful, but also offers a conceptual window into the inner mechanics of model adaptation.
@@ 5 Experimental Results
Adaptation Tasks: In the single-token setting, the output $Y_{i}$ for each sample is a single token, typically a category/class label, True/False, or MCQ choice. In the multi-token setting, the output $Y_{i}$ can be multiple tokens long and variable in length. This is the general case for open-ended generation. For both settings, we test and compare our proposed SFT pipeline on various tasks using multiple models. In many cases, we also test the adaptation methods on OOD datasets to study their behavior under distribution shift. We enlist the tasks under each category below:
@@
Data Setup: As our focus is on the data-scarce few-shot setting where ICL is typically used, we create multiple training datasets $\mathcal{D}_{\mathcal{T}}=\{(X_{i},Y_{i})\}_{i=1}^{N}$ for each task by varying $N$ in orders of 2:[2,4,8,16,]2:[2,4,8,16,...]. We create 5 different sets at each $N$ value to average out outlier effects on our final quantitative analysis. Importantly, we collect ICL activation tensors $A^{i}_{\text{ICL}}\hskip 2.0pt\forall X_{i}$ for IA2 training by using the remaining N1N-1 samples as ICL demonstrations (in random order), effectively reusing training samples. This makes sure that we use the exact same data for all adaptation methods, keeping the comparison fair. For evaluation, we sampled a different set of 500 samples for each task. Additional details about the datasets and how they are processed can be found in §A.
Multi-token details: Here, the model outputs a sequence of reasoning tokens (not inside the <<think>> tags, just standard text) before the answer in a specific format illustrated in the demonstrations. The demonstrations and expected output are both chain-of-thought (Wei et al., 2022) style. The actual answer needs to be extracted from the generated text to evaluate model performance and different datasets have different method for parsing the answer from response tokens.
  • For GSM8K and GSM8Ks, the ground truth answers have the following pattern: “<<reasoning steps>> #### <<numerical answer>>”. We do not evaluate the correctness or conciseness of the reasoning steps and only parse the numerical answer out from the response and match it exactly with the ground truth answer for calculating accuracy. We also use other common answer parsing techniques like matching with “The answer is <<value>>”, to allow for slight variations in output format. Our exact parsing function can be seen in our code.
[…]
@@
Training Setup: We use two model families to test each training method across all tasks. We use the Qwen3-4B-Base model on every task (single/multi-token). In addition, we also use the Llama-3.2-1B model on every single-token task except SciQr for which we use Llama-3.2-3B. This is because the 1B Llama model was unable to improve above random baseline for any $N$ upto 128. For the multi-token tasks, we use the Llama-3.2-1B-Instruct model as the secondary model. Using an Instruct model tests our method’s efficacy on post-trained models. For training models, we use LoRAs with rank 8 to modify the $W_{Q},W_{K}$ and $W_{O}$ matrices of self-attention layers. We chose these models and weight combinations reasonably on the basis of available resources, as we trained more than 13,000 models in total for our experiments. Training in multi-token setting is much more resource extensive. As each sample has a long answer, we decided to only use $N$ upto 16 for Llama and $N$ upto 8 for Qwen models in multi-token setting. We also fix the maximum generated tokens ($G=200$) during training for efficiency. This is used in IA2 for collecting the activation tensor $\hat{A}_{i}\in\mathbb{R}^{L\times 200\times d}\hskip 2.0pt\forall x_{i}$. For SFT, $G$ is equal to the number of tokens in the ground truth answer. Each (method, dataset) combination is trained until convergence (upto 50 epochs) with 3 learning rates (slow:1e-4, medium:3e-4, fast:1e-3), and the best performing one is chosen to neutralize the effect of hyperparameter selection on method performance. Additional details in §B.
@@ 5.2 Why is IA2 priming important?
Subspace Overlap: Although activation patterns are a footprint of the model’s functional behavior, it does not isolate changes in behavior induced by different training methods. This is because we use LoRA adaptors and the activation patterns have significant influence from the unchanged base model weights. To isolate the change in behavior, we analyze LoRA weights directly.
@@ 6 Discussion
IA2 + SFT: One natural alternative to IA2 → SFT training is IA2 + SFT, i.e., using both $\mathcal{L}_{\text{{{IA2}}}}$ and $\mathcal{L}_{\text{SFT}}$ at the same time rather than sequentially. This approach poses a practical challenge. $\hat{A}$ collected as the target activation tensor for IA2 uses the model’s own generated response tokens for any $G>1$. We call it the ICL response, which could be very different (even in length) from the ground truth response in the dataset. This makes the two objectives incompatible to train together. However, we could perform self-distillation using ICL responses, similar to Snell et al. (2022) but including an additional IA2 signal. We perform this training with a unified loss $\mathcal{L}_{\text{{{IA2}}}}+\beta\cdot\mathcal{L}_{\text{SFT}}$ and vary $\beta$ with 4 values spanning the spectrum between IA2 only and SFT only training smoothly (details in §B). After choosing the best performing (learning rate, $\beta$) combination across 5 random seeds, we report the performance for multi-token experiments in Table 3. Note that the performance of SFT only and IA2 → SFT training drops when compared to training on ground truth responses (from Table 2), but IA2 + SFT extracts significantly more performance out of ICL responses when compared to SFT only training. This highlights the richness of the IA2 training signal. Additional results of IA2 + SFT training can be found in §C.
@@
Knowledge Distillation baseline: A popular approach to distill the behavior of a stronger teacher into the student model is through soft label matching (Hinton et al., 2015). The teacher model provides a denser training signal through the probability mass associated with each token. In our work, we treat the same model (enhanced with ICL context) as the teacher. To test how IA2 based training performs in comparison to soft label matching, we train and evaluate Llama models on GSM8K (1B-Instruct) and SST2 (1B) with soft label matching on ICL responses. In Table 4, we observe that soft-label matching performs significantly better than standard SFT in the multi-token case almost reaching IA2 + SFT performance, but lacks in the single-token case. This highlights the consistency of IA2 based training in achieving high performance.
Dataset NN ICL SFT only SFT (soft labels) IA2 only IA2 $\rightarrow$ SFT IA2 + SFT
acc $\uparrow$ ece $\downarrow$ acc $\uparrow$ ece $\downarrow$ acc $\uparrow$ ece $\downarrow$ acc $\uparrow$ ece $\downarrow$ acc $\uparrow$ ece $\downarrow$ acc $\uparrow$ ece $\downarrow$
GSM8K 4 34.6 (00.9) 28.2 (04.5) 36.1 (03.5) 29.3 (05.0) 30.3 (03.2) 37.0 (01.9)
SST2 8 77.0 (02.2) 0.10 (0.02) 52.3 (04.3) 0.38 (0.14) 55.9 (08.8) 0.16 (0.06) 64.9 (15.3) 0.30 (0.05) 64.7 (19.3) 0.33 (0.19) 66.9 (18.4) 0.28 (0.21)
Table 4: Comparison with Knowledge Distillation (KD) baseline (SFT on soft labels). Although KD baseline performs better than standard SFT in the multi-token case, IA2 based training provides consistently high performance for single/multi token cases.
Why IA2 based training does not consistently beat ICL? In some multi-token cases, like the Qwen model on math datasets (GSM8K and HMathA), we find that ICL performance exceeds all training methods including IA2 based training. We suspect that this is due to mid-training of Qwen on STEM data, making ICL extremely sample efficient (surprisingly, ICL performance for GSM8K at $N=2$ is larger than that at $N=4,8$). In Table 5, we illustrate the evolution of performance with shot count. We also highlight the gap between ICL and IA2 → SFT performance, which closes quickly with more shots. Our key takeaway—IA2 → SFT performs better than SFT only—remains consistent as highlighted by the last column.
Dataset NN ICL SFT only IA2 only IA2 $\rightarrow$ SFT Gap wrt ICL Gap wrt SFT only
GSM8K 2 81.2 (00.4) 65.2 (01.0) 70.7 (01.8) 74.6 (02.4) -06.6 +09.4
4 76.4 (01.1) 70.9 (02.8) 77.4 (02.8) 73.6 (03.7) -02.8 +02.7
8 77.2 (00.9) 73.5 (01.6) 78.8 (02.5) 76.0 (02.0) -01.2 +02.5
HMathA 2 58.6 (02.0) 47.8 (04.0) 33.9 (06.4) 48.7 (07.0) -09.9 +00.9
4 60.4 (02.1) 50.4 (01.8) 47.8 (06.0) 55.3 (02.1) -05.1 +04.9
8 62.1 (01.0) 52.0 (00.8) 57.3 (03.7) 59.4 (00.8) -02.7 +07.4
Table 5: ICL is more sample efficient than training methods for Qwen models in math datasets. But IA2 $\rightarrow$ SFT quickly closes the gap in performance with increasing number of shots, and consistently beats SFT only training.
@@ 7 Related Work
Context Distillation: A large body of work targets to improve LLM adaptation using some form of compression of explicit context signals like us. These range from using intermediate contemplation tokens for compression (Cheng & Van Durme, 2024; Jiang et al., 2025), internalizing context through generating adapters on the fly (Chen et al., 2025; Charakorn et al., 2025), using related text to enhance distillation (Zhu et al., 2025; Choi et al., 2025), or compressing context directly into weights (Shen et al., 2025; Deng et al., 2024; Yu et al., 2024; Snell et al., 2022; Chen et al., 2024b; Shin et al., 2024). These works aim to distill the effect of given contexts on model outputs using token response signals. Our work is complementary to these works and improves the distillation effects through the model’s own ICL processing signals, an idea with supporting prior work (Aguilar et al., 2020; Jin et al., 2024; Yang et al., 2024).
@@ 8 Conclusion and Future Work
In this work, we used model activations to probe the functional behavior of LLMs and found a key distinction: ICL and SFT achieve adaptation through different mechanisms, and ICL often encodes richer, more generalizable patterns. Motivated by this, we introduced IA2, an SFT priming technique that enables ICL-like internal behavior in the model. This simple step shifts models into a more adaptable weight subspace, which is inaccessible to SFT alone. Extensive experiments show that IA2 consistently improves the accuracy and calibration of SFT. In the future, we aim to test other effects like impact on diversity and catastrophic forgetting of IA2 based SFT to measure how IA2 can help in post-training LLMs. We also aim to study the effect of IA2 on post-trained (RL-tuned) models.
@@ Appendix B Additional Experimental Details
LoRA details: We used LoRA rank =8=8, $\alpha=8$ and target modules are attention weights $W_{Q},W_{K}$ and $W_{O}$ for all our experiments. We settled on these parameters on the basis of generally used values for small targeted adaptations, as additional explorations require substantial compute resources for the scale of our experiments. The target modules capture the cross-attention parts ($Q,K$) and final projection $O$ to allow for better adaptability. We did preliminary ablations to make sure our results were not biased against SFT only (next paragraph). However, these selections are subject to improvement with more exploration.
[…]
FinS r8, qko 60.2 (03.0) 0.10 (0.02) 60.0 (13.6) 0.33 (0.11) 53.2 (20.4) 0.33 (0.04) 89.8 (09.0) 0.09 (0.08)
PoemS r8, qko 57.5 (01.8) 0.11 (0.02) 50.7 (03.7) 0.41 (0.07) 61.2 (12.5) 0.22 (0.06) 84.8 (08.1) 0.14 (0.08)
SST2 r8, qko 77.0 (02.2) 0.10 (0.02) 53.5 (05.0) 0.34 (0.11) 64.9 (15.3) 0.30 (0.05) 90.4 (01.3) 0.09 (0.02)
FinS r8, qk 60.2 (03.0) 0.10 (0.02) 48.8 (16.1) 0.21 (0.12) 51.9 (12.6) 0.14 (0.02) 50.4 (21.6) 0.27 (0.07)
PoemS r8, qk 57.5 (01.8) 0.11 (0.02) 48.9 (03.1) 0.24 (0.14) 42.6 (06.2) 0.13 (0.04) 56.7 (12.7) 0.29 (0.13)
SST2 r8, qk 77.0 (02.2) 0.10 (0.02) 50.0 (06.4) 0.33 (0.13) 61.6 (08.3) 0.17 (0.09) 65.7 (15.0) 0.27 (0.14)
FinS r8, qkv 60.2 (03.0) 0.10 (0.02) 54.2 (15.8) 0.34 (0.16) 66.2 (01.9) 0.16 (0.06) 87.3 (05.4) 0.06 (0.02)
PoemS r8, qkv 57.5 (01.8) 0.11 (0.02) 53.4 (05.5) 0.36 (0.11) 57.4 (08.0) 0.20 (0.08) 80.0 (07.2) 0.19 (0.07)
SST2 r8, qkv 77.0 (02.2) 0.10 (0.02) 53.5 (03.4) 0.34 (0.08) 65.3 (18.6) 0.34 (0.03) 88.7 (02.8) 0.10 (0.03)
FinS r16, qko 60.2 (03.0) 0.10 (0.02) 55.4 (13.0) 0.25 (0.12) 51.7 (19.4) 0.33 (0.04) 86.8 (09.4) 0.13 (0.09)
PoemS r16, qko 57.5 (01.8) 0.11 (0.02) 54.5 (10.9) 0.26 (0.16) 57.2 (09.1) 0.27 (0.06) 79.5 (10.0) 0.13 (0.05)
SST2 r16, qko 77.0 (02.2) 0.10 (0.02) 55.8 (13.9) 0.29 (0.14) 65.5 (14.9) 0.30 (0.05) 87.7 (05.8) 0.12 (0.06)
GSM8K r8, qko 76.4 (01.1) 70.9 (02.8) 77.4 (02.8) 73.6 (03.7)
GSM8K r16, qko 76.4 (01.1) 70.3 (05.5) 77.8 (06.4) 74.1 (03.3)
Table 7: Ablation results on LoRA parameters (rank, target modules).